OpenStudy (anonymous):

Describe the graph of the function f(x) = x3 − 11x2 + 36x − 36. Include the y-intercept, x-intercepts, and the shape of the graph.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

x-intercept is a point where line cross x-axis when y=0 so substitute y for 0 to find x-intercept y-intercept is a point where line cross y-axis when x=0 so substitute x for 0 to find y-intercept

OpenStudy (anonymous):

but i dont see any y's so how can i make it a 0?

Nnesha (nnesha):

f(x) is same as y so it's \[\huge\rm y=x^3-11x^2+36x-36\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so when x=0 y=-36

OpenStudy (anonymous):

right?

Nnesha (nnesha):

and for the shape when the highest degree is odd: if the leading coefficient is positive `Graph falls to the left and rises to the right` if the leading coefficient is negative then `graph rises to the left and falls to the right `

Nnesha (nnesha):

right

OpenStudy (anonymous):

but how do i solve for y tho? @Nnesha

Nnesha (nnesha):

you already solved for y

OpenStudy (anonymous):

not y i meant x

OpenStudy (anonymous):

cause its 0=x3 − 11x2 + 36x − 36

Nnesha (nnesha):

factor by grouping and then set it equal to zero

OpenStudy (anonymous):

(x - 2) • (x - 3) • (x - 6)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

you still there

Nnesha (nnesha):

well u can't apply the group method

Nnesha (nnesha):

how did you get that ? calculator ?

OpenStudy (anonymous):

ya

Nnesha (nnesha):

hmm alright set them equal to zero then solve for x

OpenStudy (anonymous):

its -36

Nnesha (nnesha):

that's y-intercept

Nnesha (nnesha):

(x - 2) • (x - 3) • (x - 6)=0 so x-2 =0 x-3=0 x-6=0 now solve for x

OpenStudy (anonymous):

so ad 2 6 and 3?

OpenStudy (anonymous):

x=11

Nnesha (nnesha):

don't add

Nnesha (nnesha):

there supposed to be 3 x-intercepts bec highest degree is 3

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

x=3 x=2 x=6

OpenStudy (anonymous):

is that it

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

ok what is this last part Describe the graph of the function?

OpenStudy (anonymous):

nvm we answered that already

OpenStudy (anonymous):

thank you for the help

Nnesha (nnesha):

np